What is the volume of a container that can hold 0.50 mole of gas at 25C and 1.25 atm?
1
Expert's answer
2018-03-29T07:59:41-0400
Answer:
PV = n*R*T
where:

P is the pressure of the gas (Pa), V is the volume of the gas (m3), n is the amount of substance (mole), R is the ideal, or universal, gas constant (J*mole/K), T is the absolute temperature or the gas (K).

V = n*R*T / P
V = 0.5 * 8.31 * 298 / 126656.3 = 0,00977 m3